Jenny, Katy and Bob walked 375 miles total in a walk-a-thon to raise money for the Children's Special Olympics. Jenny walked two-thirds the miles that Bob walked. Katy walked 5 more miles than Bob. How many miles did each of them walk?

Found 2 solutions by mukhopadhyay, Paul:
Answer by mukhopadhyay(490)

Let Bob walked x miles
Jenny walked 2x/3 miles
Katy walked x+5
Total distance waked by all = 375
=> x + 2x/3 + (x+5) = 375
=> 2x + 2x/3 + 5 = 375
=> (6x + 2x + 15)/3 = 375
=> 8x + 15 = 375*3 = 1125
=> 8x = 1110
=> x = 1110/8 = 555/4
Bob walked 555/4 miles, Jenny walked 555/6, and Katy walked 575/4 miles.
Verification: 555/4 + 555/6 + 575/4 = 565/2 + 555/6=(1695+555)/6 = 2250/6 = 375

Answer by Paul(988)

Bob walked x miles
Katy walked x+5 miles
Jenny walked 2x/3 miles
Equation:
2x/3+x+5+x=375
2x/3+2x=370
Multiply the whole equation by 3 to remove the fraction and simplfy:
2x+6x=1110
8x=1110
x=138.75
5+138.75=143.75
[2(138.75)]/3=92.5
Hence, Bob walked 138.75 miles, Jenny walked 92.5 miles and Katy walked 143.75 miles.
Paul.